OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if z was isolated in the first equation. 2x – 5y + z = 15 4x – y + 3z = 15 3x – 2y + 2z = 13 10x – 16y = –30 7x – 12y = –15 –2x + 14y = –30 –x + 8y = –17 22x + 16z = 90 11x + 8z = 43 –18x – 14z = –60 –5x + 8z = 43 If you were to use the substitution method to solve the following system, choose the new system of equations that would result if z was isolated in the first equation. 2x – 5y + z = 15 4x – y + 3z = 15 3x – 2y + 2z = 13 10x – 16y = –30 7x – 12y =

OpenStudy (anonymous):

Use eqn 1 therefore:\[z=15-2x+5y\]Then substitute into eqn 2:\[4x-y+3(15-2x+5y)=15\]\[4x-y+45-6x+15x=15\]\[-2x+14y=-30\] and you can already see which solution it is.